# An Efficient Intelligent Semi-Automated Warehouse Inventory Stocktaking System

### v2

@@ 3 System Overview
3.1 AI-Driven Anomaly Detection Using Isolation Forest
The proposed system utilizes Isolation Forest (IF) for anomaly detection, an algorithm well-suited for identifying outliers in high-dimensional datasets. However, other alternatives, such as DBSCAN and AutoEncoder, were considered for comparison.
@@ 3.1 AI-Driven Anomaly Detection Using Isolation Forest
Isolation Forest (IF) Efficient for high-dimensional data Sensitive to parameter tuning
Computationally efficient ($O(n\log n)$) Not optimal for time-series data
DBSCAN Detects density-based anomalies Requires manual parameter tuning ($\varepsilon$)
Robust to noise Poor performance in high dimensions
AutoEncoder (AE) Learns complex inventory patterns High computational cost
Suitable for sequential data Requires large training datasets
Why Isolation Forest?
  • It has low computational complexity and is scalable for large-scale inventory data.
  • It operates in an unsupervised manner, requiring no labeled anomaly data.
  • It is robust in handling inventory with varying SKU distributions.
  • Compared to DBSCAN and AutoEncoders, IF provides stable results without requiring extensive data preprocessing or labeled anomaly samples.
@@ 3.1.1 Data Preprocessing
Let us assume we have a dataset containing multiple handling units (HUs), where each HU contains the following information:
  • Product quantity $x_{hu}$
  • Batch $b_{hu}$
  • Bin location $bin_{hu}$
We organize the data as follows:
$\mathcal{D}=\{(x_{hu},b_{hu},bin_{hu})\},\quad hu\in\{1,2,\dots,N_{hu}\}$
@@
3.1.2 Isolation Forest Model
Isolation Forest (IForest) is a tree-based anomaly detection method typically used for high-dimensional data. In our approach, we use Isolation Forest to detect anomalies in the product quantity of each HU. Suppose we have trained an Isolation Forest model, and the model’s goal is to calculate anomaly scores based on the “isolation degree” of the sample. The anomaly score $s(x_{hu})$ is calculated as:
$s(x_{hu})=\frac{1}{T}\sum_{t=1}^{T}h_{t}(x_{hu})$
@@ 3.1.3 Batch and Bin Level Data Aggregation
    $\sigma_{b}=\sqrt{\frac{1}{N_{b}}\sum_{hu\in B_{b}}(x_{hu}-\mu_{b})^{2}}$
@@ 3.1.4 Data Fusion and Joint Feature Construction
$\mathbf{f}_{hu}=\left(x_{hu},\mu_{b_{hu}},\sigma_{b_{hu}},\mu_{bin_{hu}},\sigma_{bin_{hu}}\right)$
@@ 3.1.5 Combined Anomaly Score Calculation
For each handling unit (HU), we first compute the anomaly score based on the Isolation Forest model:
$s(x_{hu})=\frac{1}{T}\sum_{t=1}^{T}h_{t}(x_{hu})$
@@
$S_{total}(x_{hu})=\alpha\cdot s(x_{hu})+\beta\cdot\left(\frac{|\mu_{b_{hu}}-x_{hu}|}{\sigma_{b_{hu}}}\right)+\gamma\cdot\left(\frac{|\mu_{bin_{hu}}-x_{hu}|}{\sigma_{bin_{hu}}}\right)$
@@
3.1.6 Distributed Computation and Large-Scale Data Processing
@@ 3.1.7 Anomaly Detection and Real-Time Response
3.2 Enhancement of Event Ordering Using Watermark Mechanism
@@
$W(t)=\max(\text{Event Time})-\Delta t$
@@
3.2.1 Addressing Late Data Issues in Inventory Reconciliation
@@
Processed $\text{Processed Time}=\max(\text{Event Time},W(t))$
@@ 3.2.2 Handling Out-of-Order Events
In real-time systems, events often arrive out of order due to various factors such as network delays or concurrent operations at different terminals. The Watermark mechanism in Apache Flink solves this issue by providing a reliable way to handle out-of-order events. The core idea is to track the maximum event time and update the watermark to ensure that all events up to a certain point in time are processed, even if they are not received in the correct sequence.
The mathematical representation of this can be written as:
$W(t)=\max(W(t-1),\text{Event Time})\quad\text{for all event streams at time }t$
Where $W(t)$ is the updated watermark, and Event Time represents the actual timestamp of an event. The system compares the event time with the current watermark and processes the event if it is within the valid window. This mechanism ensures that out-of-order events are not discarded but rather processed correctly, maintaining inventory consistency.
3.2.3 Improving System Throughput and Real-Time Responsiveness
One of the key advantages of using the Watermark mechanism in this project is its ability to improve system throughput while maintaining real-time responsiveness. Since Flink’s Watermark mechanism ensures that events are processed based on event time, it helps maintain high throughput and low-latency processing without sacrificing data accuracy.
The throughput of the system can be mathematically modeled as:
Throughput=Number of Events ProcessedProcessing $\text{Throughput}=\frac{\text{Number of Events Processed}}{\text{Processing Time}}$
@@ 3.2.3 Improving System Throughput and Real-Time Responsiveness
The application of the Watermark mechanism significantly enhances the reliability and accuracy of the inventory management system. It ensures that inventory events are processed in the correct order and that late or out-of-order events are included in the processing sequence, maintaining data consistency and preventing reconciliation errors.
In mathematical terms, the Watermark ensures that:
Final Inventory $\text{Final Inventory State}=f(\text{All Event Times},W(t))$
@@ 3.3 ERP Synchronization and Fault Recovery
Traditional ERP systems rely on centralized databases, making them vulnerable to data loss during network failures. To enhance fault tolerance, we implemented the following mechanisms:
@@
Flink Push-Pull Synchronization Details:
  • Push Mode: Under normal conditions, Flink continuously pushes real-time inventory updates to the ERP system. Transactions are processed instantly if the ERP is operational.
  • Pull Mode: If an ERP failure occurs, Flink buffers transaction data within a durable event processing pipeline, such as Apache Kafka or RocksDB. These transactions are timestamped and stored in the Flink state backend, ensuring no data is lost.
  • Automated Recovery: Once ERP services are restored, Flink automatically pulls and resends the buffered data in the correct chronological order, ensuring consistency between warehouse transactions and ERP records.
  • Consistency Check: Before updating the ERP, Flink verifies that previously missed transactions are not duplicated, preventing inventory misalignment.
Recovery Performance:
  • Local cache storage capacity: 10,000 transactions
  • Flink event buffer size: Configurable (default: 50,000 transactions)
  • Recovery speed: Less than 1 second upon ERP reconnection
@@ 3.4 Flutter-Based Mobile Terminal
The system utilizes a Flutter-based mobile application for warehouse operators to perform stocktaking operations. The key functionalities of the mobile terminal include: - Barcode scanning for quick and efficient inventory data collection. - User-friendly interface designed for warehouse operators with minimal training requirements. - REST API communication with the backend server for real-time inventory updates.
By leveraging a cross-platform mobile application framework, the system ensures compatibility with both Android and iOS devices, making deployment more flexible in different warehouse environments.
4 Experimental Research and Analysis
4.1 Experiment Description
@@
Experimental Setup:
  • 500 inventory locations
  • 1,000 product batches
  • 37,000 storage units
  • Over 2 million individual products
Key Scenarios Tested:
  • High-density shelving: Simulates tightly packed inventory to test readability under constrained spaces.
  • Dynamic in ventory flow: Models logistics centers with high product turnover.
  • Medium to Large electronics warehouse: 20,000 square meters
4.2 Experimental Group Design
@@ 4.3 Experimental Conditions
Tag Placement and Orientation For RFID (Control Group A): RFID tags were placed uniformly across all items, with reader distances controlled (e.g., 0.5m, 1m, and 2m).
For Manual Barcode (Control Group B) and Flutter+Barcode (Experimental Group): Barcode labels were uniformly placed on items.
@@
Operator Standardization Operators across all groups were trained using a three-phase training process (see Section 4.6) to minimize performance variability.
@@
4.4 Experimental Procedure
To ensure consistency and fairness across the three groups, the experiment was conducted in three distinct phases. The procedure aimed to evaluate the speed, accuracy and usability of the inventory under controlled conditions.
@@ 4.4 Experimental Procedure
To ensure the objectivity and reliability of manual stocktaking, we implemented a double-blind review process. Two independent operators, Operator A and Operator B, conducted separate manual counts without prior knowledge of each other’s results. The final manual count was established based on the following conditions:
  • If the difference between Operator A and Operator B’s manual counts was $\leq 1\%$ , their mean value was used as the golden standard.
  • If the difference was >1%, a supervisor conducted a third verification, and the majority agreement determined the final count.
This ensured that manual data was free from individual biases and served as a reliable benchmark for evaluating the proposed system’s accuracy.
4.5 Selection and Evaluation of Experimental Participants
To ensure the reliability and consistency of the experiment, a rigorous selection process was conducted. A total of 40 candidates were initially selected based on the following criteria:
  • At least 10 years of warehouse experience
  • Familiarity with RFID inventory operations and manual recordkeeping
  • Prior stocktaking records with an accuracy of at least 95%
To further ensure competence and minimize performance variations, all candidates were required to pass a pre-test in a simulated warehouse environment, which evaluated:

### v3

@@ 3 METHODOLOGY
3.1 Anomaly Detection with Isolation Forest
The proposed system utilizes the Isolation Forest for anomaly detection, an algorithm well-suited for identifying outliers in high-dimensional datasets. However, other alternatives, such as DBSCAN and AutoEncoder, were considered for comparison, though, the algorithm comes with a few advantages compared to these alternatives. First, it has low computational complexity and is scalable for large-scale inventory data. It also operates in an unsupervised manner, requiring no labeled anomaly data. In addition, it is robust in handling inventory with varying SKU distributions, and, compared to DBSCAN and AutoEncoders, it provides stable results without requiring extensive data preprocessing or labeled anomaly samples. The comparison is summarized in Table 7.
@@ 3.1 Anomaly Detection with Isolation Forest
Isolation Forest (IF) Efficient for high-dimensional data; Computationally efficient ($O(n\log n)$) Sensitive to parameter tuning; Not optimal for time-series data
DBSCAN Detects density-based anomalies; Robust to noise Requires manual parameter tuning ($\varepsilon$); Poor performance in high dimensions
AutoEncoder (AE) Learns complex inventory patterns; Suitable for sequential data High computational cost; Requires large training datasets
@@ 3.1.1 Data Preprocessing
Let us assume we have a dataset containing multiple handling units (HUs), where each HU contains: product quantity $x_{hu}$, batch $b_{hu}$, and bin location $bin_{hu}$. We organize the data as follows:
$\mathcal{D}=\{(x_{hu},b_{hu},bin_{hu})\},\quad hu\in\{1,2,\dots,N_{hu}\},$ (3)
@@
3.1.2 Isolation Forest (IF) Model
IF is a tree-based anomaly detection method typically used for high-dimensional data. In our approach, we use the algorithm to detect anomalies in the product quantity of each HU. Suppose we have trained an IF model, and the model’s goal is to calculate anomaly scores based on the “isolation degree” of the sample. The anomaly score $s(x_{hu})$ is calculated as:
$s(x_{hu})=\frac{1}{T}\sum_{t=1}^{T}h_{t}(x_{hu}),$ (4)
@@ 3.1.3 Batch and Bin Level Data Aggregation
    $\sigma_{b}=\sqrt{\frac{1}{N_{b}}\sum_{hu\in B_{b}}(x_{hu}-\mu_{b})^{2}}.$ (6)
@@ 3.1.4 Data Fusion and Joint Feature Construction
$\mathbf{f}_{hu}=\left(x_{hu},\mu_{b_{hu}},\sigma_{b_{hu}},\mu_{bin_{hu}},\sigma_{bin_{hu}}\right),$ (8)
@@ 3.1.5 Combined Anomaly Score Calculation
For each HU, we first compute the anomaly score based on the IF model
$s(x_{hu})=\frac{1}{T}\sum_{t=1}^{T}h_{t}(x_{hu}).$ (9)
@@
$S_{total}(x_{hu})$ $=\alpha\cdot s(x_{hu})+\beta\cdot\left(\frac{|\mu_{b_{hu}}-x_{hu}|}{\sigma_{b_{hu}}}\right)$
$+\gamma\cdot\left(\frac{|\mu_{bin_{hu}}-x_{hu}|}{\sigma_{bin_{hu}}}\right),$ (10)
@@
3.1.6 Distributed Computation and Large-Scale Processing
@@ 3.1.7 Anomaly Detection and Real-Time Response
3.2 Event Ordering with Watermark
@@
$W(t)=\max(\text{Event Time})-\Delta t,$ (11)
@@
3.2.1 Late Data Issues in Inventory Reconciliation
@@
$\text{Processed Time}=\max(\text{Event Time},W(t)),$ (12)
@@ 3.2.2 Handling Out-of-Order Events
In real-time systems, events often arrive out of order due to various factors such as network delays or concurrent operations at different terminals. The Watermark mechanism in Apache Flink solves this issue by providing a reliable way to handle out-of-order events. The core idea is to track the maximum event time and update the watermark to ensure that all events up to a certain point in time are processed, even if they are not received in the correct sequence. The mathematical representation of this can be written as:
$W(t)=\max(W(t-1),\text{Event Time}),\quad\forall\text{ events at }t$ (13)
where $W(t)$ is the updated watermark, and Event Time represents the actual timestamp of an event. The system compares the event time with the current watermark and processes the event if it is within the valid window to ensure that out-of-order events are not discarded but rather processed correctly, maintaining inventory consistency.
3.2.3 Improving Throughput and Real-Time Responsiveness
One of the key advantages of using the Watermark mechanism in this project is its ability to improve system throughput while maintaining real-time responsiveness. Since Flink’s Watermark mechanism ensures that events are processed based on event time, it helps maintain high throughput and low-latency processing without sacrificing data accuracy. The throughput of the system can be mathematically modeled as:
$\text{Throughput}=\frac{\text{Number of Events Processed}}{\text{Processing Time}}.$ (14)
@@ 3.2.3 Improving Throughput and Real-Time Responsiveness
The application of the Watermark mechanism significantly enhances the reliability and accuracy of the inventory management system. It ensures that inventory events are processed in the correct order and that late or out-of-order events are included in the processing sequence, maintaining data consistency and preventing reconciliation errors. In mathematical terms, the Watermark ensures that:
$\text{Final Inventory State}=f(\text{All Event Times},W(t)),$ (15)
@@ 3.3 ERP Synchronization and Fault Recovery
Traditional ERP systems rely on centralized databases, making them vulnerable to data loss during network failures. To enhance fault tolerance, we implemented:
@@
In this Flink Push-Pull Synchronization, a few modes available. In the Push Mode, under normal conditions, Flink continuously pushes real-time inventory updates to the ERP system, hence transactions are processed instantly if the ERP is operational. In the Pull Mode, if an ERP failure occurs, Flink buffers transaction data within a durable event processing pipeline, such as Apache Kafka or RocksDB. These transactions are timestamped and stored in the Flink state backend, ensuring no data is lost. Furthermore, in the Automated Recovery, once ERP services are restored, Flink automatically pulls and resends the buffered data in the correct chronological order, ensuring consistency between warehouse transactions and ERP records. Finally, a Consistency Check is performed before updating the ERP, Flink verifies that previously missed transactions are not duplicated, preventing inventory misalignment. This setup results in a recovery performance of 10,000 transactions local cache storage capacity, default 50,000 transactions of Flink event buffer size, and recovery speed less than 1 second upon ERP reconnection.
@@ 3.4 Flutter-Based Mobile Terminal
The system utilizes a Flutter-based mobile application for warehouse operators to perform stocktaking operations. The key functionalities of the mobile terminal include barcode scanning for quick and efficient inventory data collection, user-friendly interface designed for warehouse operators with minimal training requirements, and REST API communication with the backend server for real-time inventory updates. By leveraging a cross-platform mobile application framework, the system ensures compatibility with both Android and iOS devices, making deployment more flexible in different warehouse environments.
4 EXPERIMENTS AND DISCUSSIONS
4.1 Experiment Setting
@@
The experimental setup include: 500 inventory locations, 1,000 product batches, 37,000 storage units, and over 2 million individual products. We tested several key scenarios that include: high-density shelving (simulates tightly packed inventory to test readability under constrained spaces), dynamic inventory flow (models logistics centers with high product turnover), and medium to large electronics warehouse (with 20,000 square meters space).
Three groupings are introduced to test various scenarios (see Appendix A):
@@
4.2 Experimental Procedure
To ensure consistency and fairness across the three groups, the experiment was conducted in three distinct phases, detailed in Appendix A. The procedure aimed to evaluate the speed, accuracy, and usability of the inventory under controlled conditions. Furthermore, to ensure the objectivity and reliability of manual stocktaking, we implemented a double-blind review process. Two independent operators, Operator A and Operator B, conducted separate manual counts without prior knowledge of each other’s results. The final manual count was established based on the following conditions to ensure that manual data was free from individual biases and served as a reliable benchmark for evaluating the proposed system’s accuracy:
[…]
  • If the difference was 1%, a supervisor conducted a third verification, and the majority agreement determined the final count.
4.3 Participant Selection, Evaluation, and Training
To ensure the reliability and consistency of the experiment, a rigorous selection process was conducted. A total of 40 candidates were initially selected based on the following criteria: at least 10 years of warehouse experience, familiarity with RFID inventory operations and manual recordkeeping, and prior stocktaking records with an accuracy of at least 95%. To further ensure competence and minimize performance variations, all candidates were required to pass a pre-test in a simulated warehouse environment.
Based on the results, only 13 participants who met all criteria were selected for the experiment. Furthermore, to ensure consistent performance across all groups, all operators underwent a standardized three-phase training program. This program covered the specific operations required for each experimental group and emphasized the importance of following predefined scanning protocols to ensure fairness and accuracy (see Appendix B for details).
4.4 Numerical Results and Analysis
4.4.1 Stocktaking Completion Time Analysis
The distribution of stocktaking completion time is presented in Figure 2 with key statistics summarized in Table 8. The independent samples $t$-test highlights:
  • Experimental Group vs. Group A: $t\approx-36.93$,
  • Experimental Group vs. Group B: $t\approx-106.15$, and
  • Control Group A vs. Group B: $t\approx-62.14$.
These all indicate that the corresponding $p$-values are far less than 0.0010.0010.001. In addition, the One-Way ANOVA yielded an F statistic of approximately 569656965696, demonstrating that the differences among the three groups are statistically significant ($p\ll 0.001$). Finally, for 1,000 sets of stocktaking data, the means and standard deviations for each set remain unchanged. Both the t-tests and the ANOVA confirm that the differences among the groups are extremely significant.
Figure 2: Stocktaking Completion Time Distribution
Table 8: Stocktaking Completion Time with Sample Size
Group Avg Time (s) Std (s) Samples (n)
Experimental Group 50.37 5.60 1,000
Group A 70.76 16.53 1,000
Group B 123.21 20.97 1,000
4.4.2 Inventory Accuracy and Error Analysis
To quantify the accuracy of the proposed system, we use the mean squared error (MSE), root of mean squared error (RMSE), and the 95%-confidence interval (CI) calculated as follows:
$\text{MSE}=\frac{1}{N}\sum_{i=1}^{N}(\text{System Count}_{i}-\text{Manual Count}_{i})^{2},$ (16)
$\text{RMSE}=\sqrt{\text{MSE}},$ (17)
$\text{CI}=\bar{E}\pm 1.96\times\frac{\sigma}{\sqrt{N}},$ (18)
where $\text{System Count}_{i}$ is the stock count recorded by the system, $\text{Manual Count}_{i}$ is the verified golden standard from double-blind manual stocktaking, $\bar{E}$ is the mean error, and $\sigma$ is the standard deviation of errors, and $N$ is the number of stock items tested. The results are summarized in Table 9.
4.4.3 IF Evaluation
To evaluate the IF model, we calculate the False Negative Rate (FNR) and False Positive Rate (FPR) as follows:
FNR $=\frac{\text{FN}}{\text{FN}+\text{TP}}$ (19)
FPR $=\frac{\text{FP}}{\text{FP}+\text{TN}},$ (20)
where TP (True Positive) is the actual anomalies that were correctly detected, FN (False Negative) is the actual anomalies that were missed, FP (False Positive) is the false alarms (incorrectly flagged as anomalies), and TN (True Negative) is the inventory correctly classified as normal. The summary is reported in Table 9.
Table 9: Stocktaking Performance Comparison
Inventory Accuracy IF Model
Group Accuracy MSE RMSE 95% CI TP TN FP FN FPR (%) FNR (%)
Experimental Group (Flutter + AI) 98.5% 0.85 0.92 (0.79, 0.94) 28 957 13 2 1.3% 6.7%
Group A (RFID) 97.0% 1.73 1.32 (1.61, 1.85) 60 910 20 10 2.2% 14.3%
Group B (Manual Barcode) 95.0% 2.45 1.56 (2.21, 2.68) 75 875 45 5 4.9% 6.3%
To determine the optimal parameters for the IF model, we conducted a parameter tuning experiment using Grid Search. The experiment varied the following hyperparameters: number of estimators {100, 200, 500}, max samples per tree {128, 256, 512}, contamination rate {0.005, 0.01, 0.02, 0.05}, and max features per split {3, 5, 7}. The optimized hyperparameter results are shown in Appendix C. Based on the results, we selected contamination = 0.009, as it provides the best balance between low FNR (1.3%) and FPR (1.8%), while achieving a minimal RMSE of 0.92. The results show that the AI-assisted Flutter system is faster and more accurate than both RFID and manual stocktaking methods. The use of IF for anomaly detection also provided significant improvements in anomaly detection with minimal false positives or negatives. These findings are important as they show the potential of combining AI and mobile technologies for streamlining inventory management, making it faster, more accurate, and cost-effective compared to traditional systems.
4.4.4 Participants’ Experimental Feedback
Structured interviews were conducted to gather qualitative feedback. To ensure sample diversity, participants included operators with varying experience levels and usage frequencies (from occasional to frequent). The interviews focused on five key domains, with specific questions designed for each domain. The questions for each domain and groups feedback are available in Appendix D.
4.4.5 Cost Comparison Across Scenarios
To assess the costs associated with the full RFID implementation, we estimated system acquisition and implementation expenses using data from our experimental setup. We assume that RFID readers will require periodic replacement due to technological advancements and wear-and-tear. Additionally, integration with enterprise resource planning (ERP) systems (e.g., SAP, IBM) significantly contributes to the cost.
For RFID to be fully effective, all inventory items must be tagged during implementation. If items are not tagged, inventory inaccuracy will persist, leading to increased reliance on manual stocktaking. Using 30 seconds per tag, we calculated the total labor cost required for initial tagging based on an assumed workforce wage. Finally, to determine the initial RFID tagging cost, we also estimated the average cost per tag based on a 90/10 ratio (normal vs. metal tags). During implementation, the total estimated cost of tags and labor is \$3,700. These cost breakdowns are detailed in Appendix E.
For the semi-automated barcode implementation, the costs were significantly lower than RFID. The cost of barcode tags is significantly lower than RFID. Based on an average printed barcode tag cost of \$0.04, the implementation cost for barcode tagging, including labor. The cost breakdowns for the semi-automated system are detailed in Appendix F.
We calculated the return-on-investment (ROI) period based on setup and maintenance costs, as shown in Table 10. From a cost perspective, the barcode system has significantly lower initial investment and maintenance costs, making it more suitable for businesses with limited budgets or those that do not require high-efficiency management in the short term. However, the RFID system has clear advantages in efficiency and automation, as it can quickly scan multiple items, does not require direct line-of-sight, and is resistant to contamination. It is ideal for complex supply chains and environments requiring high precision. While the initial investment for an RFID system is higher, its long-term operational efficiency and inventory management accuracy offer greater value. Overall, the RFID system is more suited for businesses that require efficient, real-time inventory management, while the barcode system is better for small businesses or organizations with limited resources that have lower efficiency demands.
Table 10: ROI Comparison
Cost Type Semi-Automated System RFID System
Initial Setup Cost \$16,830 \$545,150
Maintenance Low Medium to High
ROI Period 6–12 months 24–36 months
5 CONCLUSIONS
Next research will explore enhancements such as a combination of drone scanning with stereo camera and digital twins with federated learning for cross-warehouse predictions, and AR-assisted scanning, providing a scalable and progressive path for intelligent inventory management digitalization without disrupting existing operations. Warehouse inventory management faces persistent challenges with fully automated RFID technology, such as high implementation costs, sensitivity to environmental interference, and network load imbalances, driving the need for a more practical and cost-effective solution to enhance efficiency and accuracy in modern supply chains. This study proposes a semi-automated inventory management system that integrates barcode scanning, real-time data stream processing, and AI-driven analytics to address these issues. Experimental validation demonstrates that the system achieves a 98.5% inventory accuracy in complex warehouse scenarios, surpassing traditional RFID systems by 1.5%, while reducing equipment costs to just 3% of RFID solutions. By leveraging Apache Flink’s watermarking mechanism and Isolation Forest anomaly detection, it cuts inventory synchronization time from 45 minutes to under 5 seconds and reduces manual verification workload by 76%, offering a lightweight architecture that seamlessly connects mobile terminals, stream processing, and ERP systems. The system’s cost-effectiveness, with an 82% lower deployment cost than RFID, and its adaptability to challenging environments like metal-rich or cold chain storage, make it an ideal solution for medium to large warehouse facilities with daily turnovers below 500,000 items.
ACKNOWLEDGEMENTS
The author acknowledges the support of warehouse staff and technical teams who participated in the experimental validation of the proposed system. No specific funding sources were utilized for this research.
CONFLICTS OF INTEREST
The author declares no conflicts of interest.
DATA AVAILABILITY STATEMENTS
Due to the nature of the research, commercial supporting data is not available.
APPENDICES
Appendix A Experiment Details
The conditions are summarized in Table 11. The procedures are described in Table 12.
@@ Appendix A Experiment Details
Tag Placement and Orientation For RFID (Control Group A): RFID tags were placed uniformly across all items, with reader distances controlled (e.g., 0.5m, 1m, and 2m). For Manual Barcode (Control Group B) and Flutter+Barcode (Experimental Group): Barcode labels were uniformly placed on items.
@@
Operator Standardization Operators across all groups were trained using a three-phase training process (see Table 14) to minimize performance variability.
@@
Appendix B Pre-Test Criteria and Training Phases
The pre-test criteria for participants are summarized in Table 13. The training phases are summarized in Table 14.